HI, my name is Donesha. I want to ask a simple question about negative an positive numbers.
So say,
What would these equal:
(+)+(+)=
(+)-(+)=
(+)*(+)=
(+)/(+)=
(-)+(-)=
(-)-(-)=
(-)*(-)=
(-)/(-)=

Found 2 solutions by stanbon, jsmallt9:
Answer by stanbon(75887)

What would these equal:
(+)+(+)= +
(+)-(+)= depends : 3-4 is negative ; 4-3 is positive
(+)*(+)= +
(+)/(+)= +
(-)+(-)= -
(-)-(-)= depends : -3--4 is positive ; -4--3 is negative
(-)*(-)= +
(-)/(-)= +
==============
Cheers,
Stan H.

Answer by jsmallt9(3758)

What would these equal:
(+)+(+)= Always positive
(+)-(+)= Positive if the first number is larger, negative if the first is smaller, zero if the numbers are the same
(+)*(+)= Always positive
(+)/(+)= Always positive
(-)+(-)= Always negative
(-)-(-)= Negative if the absolute value of the first is larger, Positive if the absolute value of the first is smaller and zero if the numbers are the same
(-)*(-)= Always positive
(-)/(-)= Always positive

As you can see, the rules for subtractions are not simple. For this reason and others, I encourage everyone to learn how to change subtractions into additions. Then you only have to know how to add.